### v1

@@ Abstract
Loss of key personnel has always been a risk for research software projects. Key members of the team may have to step away due to illness or burnout, to care for a family member, from a loss of financial support, or because their career is going in a new direction. Today, though, political and financial changes are putting large numbers of researchers out of work simultaneously, potentially leaving large amounts of research software abandoned. This article presents ten tips to help researchers ensure that the software they have built will continue to be usable after they have left their present job—whether in the course of voluntary career moves or researcher mobility, but particularly in cases of involuntary departure due to political or institutional changes.
@@ Introduction
Historically, open-source research software has been longer-lived than other kinds of open-source software and government involvement has been associated with longer lifespan [9]. The challenges of responding to macro-level changes in software development support may be therefore new and unfamiliar to research software developers. In this changing environment, research software projects must be resilient. Code should outlast situations such as the author changing workplaces or careers, partnerships failing, a lab closing down, or tooling and digital infrastructure becoming unavailable. We therefore present ten tips to ensure that your software remains usable by others even if you are no longer able to work on it. Some of these tips are most useful when adopted at the beginning of a project to build resilience over time (e.g., Tips 1, 2, 5, and 7), while others are particularly relevant when a transition or disruption is imminent (e.g., Tips 3 and 8). You do not need to adopt all of them to make things better; instead, as in all emergencies, you should do what you can, where you are, with what you have.
This guide focuses on software. Other guides already exist for data [10], where archives such as Zenodo, FigShare, the End Of Term Archive or the Internet Archive are now widely used to ensure long-term availability. However, we have not found any guide to make research software resilient or to sunsetting it (i.e., phasing it out gracefully). Even if your job is not in danger, these tips will help others access and use your code, contribute to your project, and to reproduce and cite your work. As such, our recommendations are aligned with the global push for a more equitable assessment of research contributions by initiatives such as DORA, coARA, and ADORE.software.
@@
Tip 1: Stay within the law
Ensuring your work remains usable is not worth putting yourself at legal risk. Before following any of the tips below, make sure you have a legal right to do so. Most institutions and journals now have policies for licensing code and/or releasing it publicly [13, 14]. Funding agencies at various levels also often have policies (e.g., EU [^1] https://web.archive.org/web/20250424181252/https://commission.europa.eu/about/departments-and-executive-agencies/digital-services/open-source-software-strategy_en (archived 2025-04-24), UKRI [^2] https://web.archive.org/web/20250422025105/https://www.ukri.org/manage-your-award/publishing-your-research-findings/making-your-research-data-open/ (archived 2025-04-22), and NASA [^3] https://web.archive.org/web/20250422005753/https://science.nasa.gov/open-science/nasa-open-science-funding-opportunities/ (archived 2025-04-22)) which may or may not align with those of institutions and journals, and institutional policies may place restrictions on what you are allowed to say publicly about your situation. Our first actionable tip is therefore to find out what those policies are, e.g., whether you need explicit permission to publish your software, and if so, from whom.
@@
Tip 2: Define your threat model
@@
    Tip 8 explores alternatives, particularly ones that can be put in place on short notice.
@@
    Leadership threats are individual threats that affect the project’s leader, such as the leader being doxxed or personally targeted in the media because of their work. Tip 3 discusses ways to plan for this.
@@
Tip 3: Decide if you are ending, pausing, or handing off
Thousands of books describe how to start a business, but only a handful discuss handing one over or winding one down, and most of those focus on succession within family-owned businesses. Guides to running a laboratory say little more [17, 18], so such wisdom as we have is passed on person-to-person if it is passed on at all. Consider the following augmented Software Lifecycle (SLDC) shown in Figure 1, that is based on [19] and modified for this paper.
Fig 1: Augmented Software Lifecycle (SLDC) based on [19] and modified for this paper.
The first step is to decide if work on the project is ending, being suspended temporarily (for some version of “temporarily”), or if you are handing it on to someone else. If the project is ending or going on hiatus, focus on making what you have accomplished findable and citable as discussed in the tips that follow. To help you prioritize, write up the goals you initially had for this project and ask how well they were served by what you actually did. This approach could lead you to understand that the project does not have to conclude, but can be integrated into someone else’s work, even without your continued participation.
@@
However, fiscal sponsorship is only an option in some countries, and being accepted as a sponsored project is a long and challenging process. Even when it is possible, the copyright holder of the work may continue to be the institution that paid for the development of the software. Private foundations may be a more attractive alternative in some jurisdictions; as per Tip 1, check with local legal experts before taking action.
Tip 4: Choose an open license (if you can)
@@
Most discussion of open source licensing focuses on circumstances in a handful of affluent countries in the same way as discussion of fiscal sponsorship (Tip 3). For example, many Latin American countries have access to information laws that include research outcomes and artifacts. You can publish data and software under that, but the institution sets the license type, and you must use institutional repositories. As always, consult local legal experts before taking action.
Tip 5: Save everything in multiple places
Now that you are legally able to share your code, remember LOCKSS: Lots of Copies Keep Stuff Safe. But copies are not enough: the threats to your project are political as well as technological (Tip 2), so you should ensure that loss of institutional support (or worse, that institution turning on you) does not mean loss of access.
@@
When you create projects on hosted services, use a team account as the project owner rather than a personal account. Doing this makes it easier to give other people permission to manage the project, and as noted in Tip 6, a project with multiple owners from different institutions is harder for any one institution to lock down. For the same reason, do not rely solely on logins or email addresses associated with your institution: instead, ask yourself if you will still have access to the project if you lose that identity, and add an identity you personally control to the team that owns the project. You can also add your ORCID to the project’s metadata so that the project links to a profile that you control even when your contact address changes.
@@
Tip 6: Encourage community adoption
@@
Tip 7: Do what you’re supposed to (if you have time)
Research software is increasingly recognized as a citable object [30, 31, 32], and numerous books and research papers describe how to organize and run a research software project. Rather than repeating what others have said [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], this tip is a prioritized checklist of things to do if you have time. If you do not, please consult the next tip.
@@
Note that all of the suggestions above will also help with adoption, for the same reason that a tidy restaurant attracts more customers than a messy one. In addition, if you follow the suggestions above when you have time, there will be less for you to do when you don’t (as discussed in our next tip).
Tip 8: Do what you can (if you have to act quickly)
@@
    Following on from Tip 5, you can search the Software Heritage Archive with your repository url to check whether it has been saved. If your repository is not present, there is a simple process to submit it for archival.
@@
    Following on from Tip 5, you can search the Software Heritage Archive to check if your repository has been saved. If not, submitting it is a simple process.
Tip 9: Talk about what you’re doing
All of the work that we have described is extra labor. You will not want to do it, particularly not when so many other things demand attention as well. Make that work meaningful by helping others. Tell your colleagues what you have learned about your institution’s legal requirements, about navigating its bureaucracy, and about the least hurtful way to break the news to your students, collaborators, and users. Share your own tips on social media (preferably Mastodon rather than X or Bluesky, as the latter two are single points of institutional failure), and share things that haven’t worked as well so that other people can avoid the same pitfalls.

### v2

@@ Abstract
Loss of key personnel has always been a risk for research software projects. Key members of the team may have to step away due to illness or burnout, to care for a family member, from a loss of financial support, or because their career is going in a new direction. Today, though, political and financial changes are putting large numbers of researchers out of work simultaneously, potentially leaving large amounts of research software abandoned. This article presents ten rules to help researchers ensure that the software they have built will continue to be usable after they have left their present job—whether in the course of voluntary career moves or researcher mobility, but particularly in cases of involuntary departure due to political or institutional changes.
@@ Introduction
Historically, open-source research software has been longer-lived than other kinds of open-source software and government involvement has been associated with longer lifespan [9]. The challenges of responding to macro-level changes in software development support may be therefore new and unfamiliar to research software developers. In this changing environment, research software projects must be resilient. Code should outlast situations such as the author changing workplaces or careers, partnerships failing, a lab closing down, or tooling and digital infrastructure becoming unavailable. We therefore present ten rules to ensure that your software remains usable by others even if you are no longer able to work on it. Some of these rules are most useful when adopted at the beginning of a project to build resilience over time (e.g., Rules 1, 2, 5, and 7), while others are particularly relevant when a transition or disruption is imminent (e.g., Rules 3 and 8). You do not need to adopt all of them to make things better; instead, as in all emergencies, you should do what you can, where you are, with what you have.
This guide focuses on software. Other guides already exist for data [10], where archives such as Zenodo, FigShare, the End Of Term Archive or the Internet Archive are now widely used to ensure long-term availability. However, we have not found any guide to make research software resilient or to sunsetting it (i.e., phasing it out gracefully). Even if your job is not in danger, these rules will help others access and use your code, contribute to your project, and to reproduce and cite your work. As such, our recommendations are aligned with the global push for a more equitable assessment of research contributions by initiatives such as DORA, coARA, and ADORE.software.
@@
Rule 1: Stay within the law
Ensuring your work remains usable is not worth putting yourself at legal risk. Before following any of the rules below, make sure you have a legal right to do so. Most institutions and journals now have policies for licensing code and/or releasing it publicly [13, 14]. Funding agencies at various levels also often have policies (e.g., EU [^1] https://web.archive.org/web/20250424181252/https://commission.europa.eu/about/departments-and-executive-agencies/digital-services/open-source-software-strategy_en (archived 2025-04-24), UKRI [^2] https://web.archive.org/web/20250422025105/https://www.ukri.org/manage-your-award/publishing-your-research-findings/making-your-research-data-open/ (archived 2025-04-22), and NASA [^3] https://web.archive.org/web/20250422005753/https://science.nasa.gov/open-science/nasa-open-science-funding-opportunities/ (archived 2025-04-22)) which may or may not align with those of institutions and journals, and institutional policies may place restrictions on what you are allowed to say publicly about your situation. Our first actionable rule is therefore to find out what those policies are, e.g., whether you need explicit permission to publish your software, and if so, from whom.
@@
Rule 2: Define your threat model
@@
    Rule 8 explores alternatives, particularly ones that can be put in place on short notice.
@@
    Leadership threats are individual threats that affect the project’s leader, such as the leader being doxxed or personally targeted in the media because of their work. Rule 3 discusses ways to plan for this.
@@
Rule 3: Decide if you are ending, pausing, or handing off
Thousands of books describe how to start a business, but only a handful discuss handing one over or winding one down, and most of those focus on succession within family-owned businesses. Guides to running a laboratory say little more [17, 18], so such wisdom as we have is passed on person-to-person if it is passed on at all. Consider the following augmented Software Lifecycle (SLDC) in Fig. 1, that is based on [19] and modified for this paper.
Fig 1: The software lifecycle.
The first step is to decide if work on the project is ending, being suspended temporarily (for some version of “temporarily”), or if you are handing it on to someone else. If the project is ending or going on hiatus, focus on making what you have accomplished findable and citable as discussed in the rules that follow. To help you prioritize, write up the goals you initially had for this project and ask how well they were served by what you actually did. This approach could lead you to understand that the project does not have to conclude, but can be integrated into someone else’s work, even without your continued participation.
@@
However, fiscal sponsorship is only an option in some countries, and being accepted as a sponsored project is a long and challenging process. Even when it is possible, the copyright holder of the work may continue to be the institution that paid for the development of the software. Private foundations may be a more attractive alternative in some jurisdictions; as per Rule 1, check with local legal experts before taking action.
Rule 4: Choose an open license (if you can)
@@
Most discussion of open source licensing focuses on circumstances in a handful of affluent countries in the same way as discussion of fiscal sponsorship (Rule 3). For example, many Latin American countries have access to information laws that include research outcomes and artifacts. You can publish data and software under that, but the institution sets the license type, and you must use institutional repositories. As always, consult local legal experts before taking action.
Rule 5: Save everything in multiple places
Now that you are legally able to share your code, remember LOCKSS: Lots of Copies Keep Stuff Safe. But copies are not enough: the threats to your project are political as well as technological (Rule 2), so you should ensure that loss of institutional support (or worse, that institution turning on you) does not mean loss of access.
@@
When you create projects on hosted services, use a team account as the project owner rather than a personal account. Doing this makes it easier to give other people permission to manage the project, and as noted in Rule 6, a project with multiple owners from different institutions is harder for any one institution to lock down. For the same reason, do not rely solely on logins or email addresses associated with your institution: instead, ask yourself if you will still have access to the project if you lose that identity, and add an identity you personally control to the team that owns the project. You can also add your ORCID to the project’s metadata so that the project links to a profile that you control even when your contact address changes.
@@
Rule 6: Encourage community adoption
@@
Rule 7: Do what you’re supposed to (if you have time)
Research software is increasingly recognized as a citable object [30, 31, 32], and numerous books and research papers describe how to organize and run a research software project. Rather than repeating what others have said [33, 34], this rule is a prioritized checklist of things to do if you have time. If you do not, please consult the next rule.
@@
Note that all of the suggestions above will also help with adoption, for the same reason that a tidy restaurant attracts more customers than a messy one. In addition, if you follow the suggestions above when you have time, there will be less for you to do when you don’t (as discussed in our next rule).
Rule 8: Do what you can (if you have to act quickly)
@@
    Following on from Rule 5, you can search the Software Heritage Archive with your repository url to check whether it has been saved. If your repository is not present, there is a simple process to submit it for archival.
@@
    Following on from Rule 5, you can search the Software Heritage Archive to check if your repository has been saved. If not, submitting it is a simple process.
Rule 9: Talk about what you’re doing
All of the work that we have described is extra labor. You will not want to do it, particularly not when so many other things demand attention as well. Make that work meaningful by helping others. Tell your colleagues what you have learned about your institution’s legal requirements, about navigating its bureaucracy, and about the least hurtful way to break the news to your students, collaborators, and users. Share your own rules on social media (preferably Mastodon rather than X or Bluesky, as the latter two are single points of institutional failure), and share things that haven’t worked as well so that other people can avoid the same pitfalls.